• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name paschke

    Log entry time 00:02:21 on May24,2004

    Entry number 122570

    keyword=Swing shift summary






    Shift starts with beam returning. Harp scans on previous shift
    showed a large spot (390um x370 um at harp 6 and about 170 um at
    hapr 5, so increasing in size toward the target). It was decided
    that the spot size issue will be approached after the optics had been
    adjusted for septum steering.

    16:00 beam here, spot on BeO target shows irregular, occasional size
    and position jitter. MCC says this is result of a "wiggle' in the
    injector Reza Kazimi is fixing it.

    16:18 MCC says they're taking beam away to fix "synch light / tune mode
    problem". I understood the part about taking beam away...

    16:35 Beam returns. MCC is working on FFB. I get them to put the OTR in and
    change filters until data is valid. Spot size is reasonable, with
    OTR width of < .2 mm suggests <5e-5 E spread.

    17:05 Beam off to start ramping septum. Moved to Empty target, as MCC
    said that the BeO would make it more difficult to view the beam at
    the dump.

    18:00 Right Septum. Set I = 231.45 A, Read I = 231.25 A,
    Temp on coils went from about 7.5K/7K to 6.7K/6.25K
    Left Septum. SetI = 230.52 A, Read I = 230.25 A
    Temp on coils stayed roughly constant, about 6.4K/6.2K
    Beam returns to steer to dump

    18:10 Septa on, beam on dump diffuser.
    BPM readings: 4a(x=2.065,y=0.07), 4b(x=-1.927,y=0.06)
    Beam off to insert BeO viewer.

    18:17 Ready to come back, but MCC will take beam away for a few
    minutes for injector work to clean up tail on Hall B beam.

    18:30 Beam returns. Starting harp scan.
    Harp scan 1961, Coda run 1061.
    series of coda runs while MCC considers which quads to change. Runs 1063-1066
    Harp scan 1962, Coda run 1067
    Harp scan 1963 Spot at 3B is about 250um x 250 um. This wasn't
    converging quickly, so I'll consult with spokespeople before spending
    more time.

    19:10 turned on raster, started checking spot size.
    Run 1068 got 1.8mm x 1.5mm. (Mag currents 10.5 A x 10.6 A)
    Run 1069. Requested 2.2x2.5, got 2mm x 2mm (Mag currents 11.6 A x 13.2A)

    Run 1070, Changed prescales to emphasize coincidence data.
    19:25-19:35 DAQ hang. Run 1071 started ok.

    Analyzer says that spot size is 1.75mm x 2mm. Requested 2.7x2.5mm.
    (Mag currents 14.22 A x 13.24 A)
    Run 1072, raster requested 2.7mm x 2.5mm

    19:47 Moved to 15 cm LH2 target. raster on.
    Run 1073 First Data on LH2! 5uA. No visible change in target parameters,
    small change in septum temperature.
    Changed prescale to match suggested trigger rates at 25 uA.

    Run 1074 new prescale settings on LH2. At 13 uA, we see 900 Hz in DAQ, so
    I mis-calculated. RICH tripped at the this current, but recovered
    when reset without lowering beam current.
    One coil on Right Septum reached 7.2 K.

    Run 1075 Reduced current to 9 uA. (LH2)

    Lowered RICH HV from 2050 V to 2000 V

    Run 1076 RICH HV = 2000 V. 9 uA. DAQ rate about 450 Hz.
    Run 1077 RICH HV = 2000 V. Short run, stopped to switch target

    21:20 Moved to 4 cm target.

    Run 1078 Current ramped to 30 uA

    21:38 Noticed target stripcharts have gone flat. Rebooted target IOC.

    Run 1079 4 cm LH2. 30 uA.
    Run 1080 4 cm LH2. up to 35 uA.

    22:39 Moved to target labeled "10 cm Optics" (which is actually 24 cm optics target)
    Run 1081 (bad prescale, try again)
    Run 1082 10 uA, raster on. Data shows this is actually the 24 cm Optics target
    Run 1083 10 uA, raster off.

    22:55 Moved to target labeled "24 cm Optics" (actually 10 cm target)

    Run 1084 10 uA, raster off.
    Run 1085 10 uA, raster on.

    23:10 Moved back to 4 cm LH2

    Run 1086 Current to 40 uA

    23:35 Moved to 5 uA. bpm 4b reads back uniformly zero.
    23:40 BPM 4b still not working. Turned beam off for pedestal run.

    Run 1087 Pedestal run

    23:45 pedestal run over, but BPM 4b still not working.
    23:50 IOC reboot didn't help, EOC paged for help on BPM 4b.